OpenStudy (anonymous):

Please and Thank you! "Evaluate each of the following. In each case, explain your thinking."

OpenStudy (anonymous):

I'm not looking for the answers, just some guidance. Besides I did a few already...

OpenStudy (anonymous):

a. Part I: √3/2 b. Part I: 5 = c Part II: don't I just use the equation given to me?\[\sin (\tan^{-1} (\frac{ 4 }{ 3 }))\]

OpenStudy (psymon):

Use the triangle given to you. What the question is asking is what does sin x = using a triangle where tanx = 4/3? So using that triangle, just find the sin of it :3

OpenStudy (anonymous):

OH!!! gotcha \[\sin = \frac{ opp }{ hyp }\]

OpenStudy (psymon):

Yeah xD

OpenStudy (anonymous):

\[\sin = \frac{ 4 }{ 5 }\] Is that my answer?

OpenStudy (psymon):

Yep.

OpenStudy (anonymous):

Final answer?

OpenStudy (psymon):

Well, it wants the hypotenuse, so that was 5. Now we have sin(theta) = 4/5. you probably should actually find theta from the looks of the question

OpenStudy (psymon):

Geez, amazing how many of the people asking questions on here are all people using this apex whatever.

OpenStudy (anonymous):

Yes! It is crazy :P

OpenStudy (psymon):

Some questions ive seen 4 times I think, lol. So you dont still use apex, do ya?

OpenStudy (anonymous):

Not exactly? I used to though. I can still log on, and check out the different assignments. I want to be ready for my first class "Intro to College Math"

OpenStudy (anonymous):

So asking people about certain formulas and identities should help! Also I got:\[\theta = 53.13º\]

OpenStudy (psymon):

Lol, that works xD And that seems right to me. But yeah, intro to college math? I would think trig is above the intro level xD

OpenStudy (anonymous):

Oh. So then I was doing all of this for NO REASON!?!?!?! NOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOO

OpenStudy (anonymous):

But actually, its okay. I still need to keep my mind sharp!

OpenStudy (psymon):

Well......At my college, we have math 124 (college algebra). Which is at the same level ish as math 126 (precalculus 1). Then we have math 127 (precalculus 2/trigonometry). So at my school, intro to college math would be 1-2 levels below trig.

OpenStudy (anonymous):

oh okay! Well, I only have to take one course for Math (for my major)

OpenStudy (psymon):

Ah, nice, lol. I still have 4 more x_x It sucks, if it werent for a schedule conflict id be able to take calc 3 this semester.

OpenStudy (anonymous):

Oh dang! I don't think I'd ever be able to do that.. What is your major?

OpenStudy (psymon):

Well, I have an associates in japanese and im working on a math major. Im attending a 4-year uni for the first time starting the 27th where Im taking more Japanese classes for a minor and starting on my bachelors in math. I still have 5 more classes at the community college level first, though, but since I have a high enough math already, I can start taking some 300 level classes.

OpenStudy (anonymous):

JAPANESE? I bow down man! And 300 level? I'm still at 100, maybe less... So inadequate

OpenStudy (psymon):

まだ日本語を学んでいるですが、もうたくさん習いました

OpenStudy (psymon):

Eh, my japanese isnt that great, but I know some, haha. But yeah, I guess. All 300 level means is Ive done enough calculus, haha.

OpenStudy (anonymous):

I have no clue what that is... I've taken elementary Spanish, one year of high school Mandarin, and one year of Tagalog/Ilocano

OpenStudy (anonymous):

Still know NOTHING! :P

OpenStudy (psymon):

Lol, well after 4 semesters of Japanese I still feel like I know nothing xD My hearing is horrible, so I have no idea what people are sayign to me. but I know how to occasionally answer and such. I can't speak well either, haha. But I'm better than I thought I'd be if Im just typing or texting in Japanese, lol.

OpenStudy (anonymous):

Why is your name Psymon? Psy is a famous rapper from South Korea, but you probs knew that! And it sounds like "simon"

OpenStudy (anonymous):

Sorry, I was just always wondering!

OpenStudy (psymon):

Yeah, it does sound like Simon, lol. I've had that pseudonym thingy since I was maybe 10 or 11, lol. So something that has stuck. If I don't use Psymon, I use PsymonSays

OpenStudy (anonymous):

I see! Oh and we kinda never finished the problem? Letter a, Part II?

OpenStudy (psymon):

To solve the problem, find the the angle whose value is rt3/2? That problem?

OpenStudy (anonymous):

Wait... what?

OpenStudy (psymon):

Im not exactly sure which question xD I see the arccos(cos-pi/6) one

OpenStudy (anonymous):

...find the angle in the interval [0,π] whose cosine is √3/2? Yes that one. Sorry, I was a bit distracted

OpenStudy (psymon):

Well, easy if you know your unit circle xD It gives you the answer with the whole arccos(cos-pi/6). Since arccos and cos are inverses, they just cancel out and leave -pi/6. Only thing to know is that since cosx is an even function, cos(-pi/6) is the same as cospi/6.

OpenStudy (anonymous):

Doesn't cos(-π/6) already equal √3/2?

OpenStudy (psymon):

Yep. But it wants an angle between 0 and pi xD

OpenStudy (anonymous):

30º angle!

OpenStudy (psymon):

Lol, yep xD But all I meant is that any cosine angle that is negative is the same value when its positive. cos(-pi/6) = cos(pi/6)

OpenStudy (anonymous):

I understand that part... And both equal 30º.